Crew 213 Operations Report 24-May-2019

Crew 213 Operations Report 24-May-2019

SOL: 6

Name of person filing report:Dana Levin

Non-nominal systems: None

Generator (hours run): overnight 2000-0700

Solar— SOC % (Before generator is run at night): 80% (see above)

Diesel Reading – 73%

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons.

Water (auxillary tank) – 0

Water (static tank) – 400 gallons

Auxillary to Static tank transfer– Yes

Gallons transferred: 100

Water in GreenHab – Greenhab ops have closed for the season, Crew 213 will not be performing greenhab ops

Water (loft) – Static to Loft Pump used – yes

Water Meter: 144250

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Personal rover of director used with her permission by CU- Staff

Hours: 1 hour

Beginning charge: 100%

Ending charge: 84

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 80

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 79

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 82

Currently charging: yes

Notes on rovers: nothing to report

ATV’s Used: no

Reason for use: N/A

Oil Added? no

ATV Fuel Used: no

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs are safely secured near directors residence and will not be used by crew 213 for planned nominal operations

HabCar used and why, where? Yes, to obtain water from town.

CrewCar used and why, where? no (in Grand Junction)

General notes and comments:

– Crew 213 had successful simulation of several medical contingencies on EVA today.

Summary of internet: functional

Summary of suits and radios: Suits are all working well. We are using our own radios.

Summary of Hab operations: nominal

Summary of GreenHab operations: none

Summary of ScienceDome operations: Used for course lectures and briefings. All systems nominal

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: None at this time

Questions, concerns and requests to Mission Support: None needed. Thank you so much for your support of our mission.

Crew 213 Mission Summary 24-May-2019

Crew 213 Mission Summary 24-May-2019

This course was our largest and most extensive iteration of our combined medicine and engineering education program. It went extremely well. Our crew enjoyed the simulations and clearly learned a lot about medicine and engineering as it relates to spaceflight.We continue to be grateful to the Mars Society for the opportunity to use this facility and all the resources it offers in our educational efforts. The major challenges we encountered were unpredictability with weather and managing a larger than normal number of students. The media team was an asset to our course and assisted us with arranging lighting and effects for our simulations.

As has been typical of our missions the daily EVA scenarios were handled safely and effectively and the emergency simulations were coupled with debriefs to ensure effective transfer of each learning objective. The feedback received from the crew both informally and through our own feedback process indicated a high level of enjoyment, respect for the facility, the course, and the challenges of a mission to Mars.

Our experience teaching doctors in previous versions of this course have lead us to a robust didactic, simulation, and discussion based curriculum. The landscape and the difficulties of living in the habitat are well known to us so there were few additional surprises. However, the weather was uncommonly rainy which forced us to amend our day to day operations on a near continuous basis. In the future, we will plan for back up activities to keep students engaged even when the weather precludes outside activities.

As this is the third time we’ve incorporated research projects into our curriculum, it has become a relatively routine part of our instruction. The crew was very receptive as were outside parties and we are looking forward to expanding this work in future missions. Our research is primarily focused on habitability, rapid iterative design, and feedback from task saturated personnel. We hope to present this at future meetings and continue to solicit more projects that can benefit from our unique population of medical professionals. Our projects for this year included a training and testing session of just in time training for ultrasound guided nerve blocks.

As always the realism of the EVA landscape is the most impressive feature of the MDRS site. The habitat facilities, EVA suits, and food supplies are well suited to the experience, however we have noticed a need for maintenance in both the habitat and space suits. Thank you for the continued opportunity to work with you on this project, we look forward to our continued collaboration.

Operations Report- May 12th

Crew 212 Operations Report 12-MAY-2019

SOL: 7

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): 11:15

On (last night): 8:15 PM

Off (this morning): 7:30 AM

Solar— SOC 61% at 7:45 PM [From Director, a
check will be run earlier in the afternoon from now on, to
avoid crossing the 80% limit again).

 

Diesel Reading – 90 %

 

Propane Reading – 46% [correcting the 45% of
yesterday]

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 100 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 140 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143547.8

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.8h

Beginning charge:

Ending charge: 79%

Currently charging: Yes

Opportunity rover used: No

Hours: 59.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: 95.1h

Beginning charge: 100%

Ending charge: 78%

Currently charging: Yes

Notes on rovers: Curiosity was taken for the Engineering EVA that was performed to free the mouse (see general notes and comments)

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: During the night, a small mouse was captured with one of the two traps in the Hab. The morning, after a scheduled EVA was concluded, an Engineering EVA was organised to free the mouse close to Cowboy’s Corner. Crew Engineer Paolo and Crew Journalist Zoe travelled for about 20 minutes with Curi to the location and freed the mouse in a bush. On the way back there was a small problem with the radios, but it was solved as it was determined that a button was incidentally pushed while mounting the rover. Today was a resting day for most of the crew, and we had birthday cookies for our Spacesuit engineer Camilo.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: A mouse was found on the upper deck around 1:45AM; it was inside one of the rooms and then moved to another, passing under the doors. A trap was placed and the mouse walked into it. An Engineering EVA was performed in the morning to release it. Other points: it was not possible to fix the oven; the Space chess board is missing a pawn; the leak from the pump has now penetrated the roof in the commander’s room through the loft floor plywood.

Summary of GreenHab operations: some sage was collected for lunch. A lizard was found inside, but it escaped once it saw a crew member.

Summary of ScienceDome operations: cleaning was performed by the crew.

Summary of RAM operations: Drone cleaning, repairs and testing were performed.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: we need a static water tank refill. We have run out of kitchen paper.

Crew 197 arrives at the MDRS in 10 days!

Crew 197 (VR CrowdExplorers)

Overall Goals of Mission

  1. Create a research whitepaper for using VR to train analog simulation participants.
  2. Pilot out MarsVR Training scenarios and script/enact new scenarios.
  3. Update interior/exterior scans for MarsVR future releases.
  4. Capture video & 360 footage for future promotional videos.
  5. Prepare the station for beginning of field season.
    • Installation of new Donor & Memorial plaques. (James)
    • Start up some plants in the GreenHab.
    • Setup of new HAL workstation & displays. (James)
    • Updates to Internet Router firmware & configuration. (James)
    • Other logistical tasks.

Crew will be IN SIM starting the afternoon of Sol 1 (October 22).

MDRS Photogrammetry Scans Completed

The MarsVR team has completed their scans of the MDRS facilities and terrain.  Several Matterport 360 walkthoughs are already available, and the team will be working this summer to create the full photo-realistic VR environment for MDRS crew training, that will be put into use for next field season.

Read More

Call for help on VR project – June 19th – 25th

Hi all,
To support the creation of a VR environment for future MDRS crew training and research, we will be conducting scans and drone operations at the MDRS starting on June 19th for around a week.  We’d like to get help from a past crewmember at the MDRS.  Please let us know if you are interested in helping out with our project – we can work out your travel arrangements.  Please forward this to anybody that might be interested in participating.
More info on our project is at: http://marsvr.io

Thanks,

James Burk

Program Manager, MarsVR

marsvr@marssociety.org